# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v1

@@ A.8 Hyper-Parameter Study with Respect to Target Separation Margin
Tables 16 and 17 consider varying the target margin $\tau^{mm}$ when learning EDDP. For EfficientNet, increasing the target classification margin (i.e. the linear separability of patch embeddings) leads to a substantial improvement in both $\mathcal{S}^{1}$ and $\mathcal{S}^{2}$ for both EDDP-C and EDDP-U. We stress the fact that increasing the margin is accomplished by decreasing $\tau^{mm}$, as the margin is inversely proportional to the target loss. For ResNet18, the results are more mixed: for EDDP-U, decreasing the margin improves $\mathcal{S}^{1}$ and $\mathcal{S}^{2}$, while for EDDP-C there is an improvement in $\mathcal{S}^{1}$ but at the cost of a slightly inferior $\mathcal{S}^{2}$ score. We emphasize that when $\tau^{mm}=4.0$, the results for EfficientNet are substantially better than the metrics we reported in the main body of the paper, for which we used $\tau^{mm}=5.0$ (Section A.7).
A.9 Additional Experiments on the Faithfulness Assessment of the Encoding-Decoding Direction Pairs
Figures 27 and 28 complement the experiments presented in Figures 11 and 12 for the remaining combinations regarding $\mathcal{L}^{fso}$. Even though less prominent in ResNet18, together with Section 7.2 the experiments sufficiently support that $\mathcal{L}^{fso}$ pushes the distribution of cosine similarities between dreaming directions and signal vectors to become more bell-shaped. Overall, the experiments in this section, similar to the experiments presented in Section 7.2, indicate that in approximately 90% of the cases, these cosine similarities exceed 0.7. In a less rigorous but more qualitative sense, we found that when learning EDDP with $\mathcal{L}^{fso}$, this loss converges to values close to $\approx 0.02$, whereas measuring the loss value for a direction set learned without $\mathcal{L}^{fso}$ often leads to a value close to $\approx 0.1$, which is still not that far from indicating orthogonality between filters and signal vectors.
A.10 Ablation Study with Respect to Filter-Signal Orthogonality Loss
@@
In this Section we conduct a detailed ablation study with respect to the use of Filter-Signal Orthogonality Loss $(\mathcal{L}^{fso})$ that was introduced in Section 4.3. In Section A.9 we discussed the impact of $\mathcal{L}^{fso}$ on the faithfulness of the direction pairs, while in this Section we discuss the effects of $\mathcal{L}^{fso}$, in interpretability and influence terms. Tables 18, 19, 20, 21, 22, 23 depict results in terms of interpretability and influence metrics.
@@
A.11 Comparison with Unsupervised Interpretable Basis Extraction and Concept-Basis Extraction in Practical Experiments
Table 24: Comparison and Ablation with respect to method enhancements that we propose in this work. $\mathcal{L}^{al}$ stands for Augmented Lagrangian Loss and $\mathcal{L}^{cc}$ refers to the CNN Classifier loss that was proposed in Doumanoglou et al. (2024). Methods below the dashed line correspond to variations that use the contributions we make in this work. Here the network is ResNet18 and $I=512$.
@@
Table 25: Comparison and Ablation with respect to method enhancements that we propose in this work. $\mathcal{L}^{al}$ stands for Augmented Lagrangian Loss and $\mathcal{L}^{cc}$ refers to the CNN Classifier loss that was proposed in Doumanoglou et al. (2024). Methods below the dashed line correspond to variations that use the contributions we make in this work. Here the network is ResNet50 and $I=2048$.
@@
Tables 24 and 25 briefly present comparisons of our work with the previous works that we extend, in terms of interpretability. First, with other aspects of Concept Basis Extraction (CBE) (Doumanoglou et al., 2024) being intact, we assess the efficacy of our Unconstrained Uncertainty Region Alignment loss in improving the interpretability of the clustering, compared to the CNN Classifier Loss $\mathcal{L}^{cc}$ that was proposed previously in Doumanoglou et al. (2024). This case is referred to in the tables as CBE /w $\mathcal{L}^{uur}$. We observe that in three out of the four cases, our Uncertainty Region Alignment loss leads to a notable relative improvement, by up to 22.56% in $\mathcal{S}^{2}$, which indicates a clustering with improved concept diversity. When we additionally take into account the rest of our contributions, and compare EDDP-U with CBE, we find that in the same three out of four cases, interpretability is further improved by up to 58.8% in terms of $\mathcal{S}^{1}$. Transitioning from CBE to EDDP-C, may lead to further improvements in $\mathcal{S}^{2}$ by up to 33.5%, yet with a decreased score in terms of $\mathcal{S}^{1}$.
A.12 More Qualitative Segmentations and Statistics for Evaluating the Interpretability of the Concept Detectors
Figures 29, 30, 31, 32, 33, 34, 37, 38, 39 depict qualitative segmentation results obtained using the concept detectors that were learned with the proposed method for the architectures that we studied in this work. Visualizations are obtained using Bau et al. (2017). We can verify that concept detectors appear to be monosemantic. In some cases, there are more than one concept detector detecting the same concept. However, in most cases, the sets of positively classified samples across detectors are disjoint. Figures 35 and 36 depict qualitative segmentations obtained via NMF and PCA. In some of the depicted examples, it is evident that the identified concepts are less monosemantic. Tables 26, 27, 28 and 29 summarize the statistics obtained by Network Dissection regarding the interpretability of the clusterings considered in this work.
@@ II Method Color Object Part Material Action Scene Texture Total
A.13 More Global Model Explanations via Concept Sensitivity Testing
@@
A.14 More Local Explanations with Concept Contribution Maps
@@
A.15 Relation to Sparse Auto-Encoders
The superposition hypothesis Elhage et al. (2022) assumes that neural networks linearly represent more features than there are neurons in their hidden layers. In this line of work, a feature is defined to be an abstract property of the input, which may or may not align with human intuition, and is exploited by the network to make predictions. Sparse AutoEncoders (SAEs) Bricken et al. (2023) have been proposed as a tool to take features out of superposition; that is, given the activations of a network’s intermediate layer, SAEs try to linearly decompose these activations in terms of latent feature components under a sparsity objective.
Let ${\bm{x}}\in{\mathbb{R}}^{D}$ denote the activation of a network in a hidden layer of study and $I\gg D$ denote the number of latent features that the network is assumed to represent. While several SAE variants have been proposed, such as Lim et al. (2024); Cunningham et al. (2024); Sharkey et al. (2022); Bussmann et al. (2025), in their baseline form Bricken et al. (2023) they first extract latent feature components $\bm{v}$} by:
@@ A.15 Relation to Sparse Auto-Encoders
Despite the conceptual similarities between our method and SAEs, we also have several differences. First, (41) is conceptually acting as a classifier with positive predictions whenever the extracted signal value exceeds the threshold in ${\bm{b}}_{\text{enc}}$. From one perspective, the same equation acts like centering the signal value around the classification threshold for the presence of the concept in the representation. In SAEs, sparsity is enforced in the units of the latent features $\bm{v}$}, while in our approach, the ReLU activation function is replaced with a sigmoid, and thus sparsity is enforced in the semantic space of concepts. Second, in SAEs, ${\bm{W}}_{\text{dec}}$ are learned via a reconstruction objective, while in our approach, signal vectors are learned under the properties of a probabilistic model constrained on the extracted signal values and the feature activations themselves. Third, we also consider constraining each decoding direction to be perpendicular to the signal vectors of other concepts, for exact signal value extraction. Finally, we also consider uncertainty region alignment which additionally exploits the use of the directions by the model.

### v2

@@ A.13 Hyper-Parameter Study with Respect to Target Separation Margin
Tables 17 and 18 consider varying the target margin $\tau^{mm}$ when learning EDDP. For EfficientNet, increasing the target classification margin (i.e. the linear separability of patch embeddings) leads to a substantial improvement in both $\mathcal{S}^{1}$ and $\mathcal{S}^{2}$ for both EDDP-C and EDDP-U. We stress the fact that increasing the margin is accomplished by decreasing $\tau^{mm}$, as the margin is inversely proportional to the target loss. For ResNet18, the results are more mixed: for EDDP-U, decreasing the margin improves $\mathcal{S}^{1}$ and $\mathcal{S}^{2}$, while for EDDP-C there is an improvement in $\mathcal{S}^{1}$ but at the cost of a slightly inferior $\mathcal{S}^{2}$ score. We emphasize that when $\tau^{mm}=4.0$, the results for EfficientNet are substantially better than the metrics we reported in the main body of the paper, for which we used $\tau^{mm}=5.0$ (Section A.12).
A.14 Ablation Study with Respect to Filter-Signal Orthogonality Loss
@@
In this Section we conduct a detailed ablation study with respect to the use of Filter-Signal Orthogonality Loss $(\mathcal{L}^{fso})$ that was introduced in Section 4.3. In Section A.11 we discussed the impact of $\mathcal{L}^{fso}$ on the faithfulness of the direction pairs, while in this Section we discuss the effects of $\mathcal{L}^{fso}$, in interpretability and influence terms. Tables 19, 20, 21, 22, 23, 24 depict results in terms of interpretability and influence metrics.
@@
A.15 Comparison with Unsupervised Interpretable Basis Extraction and Concept-Basis Extraction in Practical Experiments
Table 25: Comparison and Ablation with respect to method enhancements that we propose in this work. $\mathcal{L}^{al}$ stands for Augmented Lagrangian Loss and $\mathcal{L}^{cc}$ refers to the CNN Classifier loss that was proposed in CBE. Methods below the dashed line correspond to variations that use the contributions we make in this work. Here the network is ResNet18 and $I=512$.
@@
Table 26: Comparison and Ablation with respect to method enhancements that we propose in this work. $\mathcal{L}^{al}$ stands for Augmented Lagrangian Loss and $\mathcal{L}^{cc}$ refers to the CNN Classifier loss that was proposed in CBE. Methods below the dashed line correspond to variations that use the contributions we make in this work. Here the network is ResNet50 and $I=2048$.
@@
Tables 25 and 26 briefly present comparisons of our work with the previous works that we extend, in terms of interpretability. First, with other aspects of Concept Basis Extraction (CBE) (CBE) being intact, we assess the efficacy of our Unconstrained Uncertainty Region Alignment loss in improving the interpretability of the clustering, compared to the CNN Classifier Loss $\mathcal{L}^{cc}$ that was proposed previously in CBE. This case is referred to in the tables as CBE /w $\mathcal{L}^{uur}$. We observe that in three out of the four cases, our Uncertainty Region Alignment loss leads to a notable relative improvement, by up to 22.56% in $\mathcal{S}^{2}$, which indicates a clustering with improved concept diversity. When we additionally take into account the rest of our contributions, and compare EDDP-U with CBE, we find that in the same three out of four cases, interpretability is further improved by up to 58.8% in terms of $\mathcal{S}^{1}$. Transitioning from CBE to EDDP-C, may lead to further improvements in $\mathcal{S}^{2}$ by up to 33.5%, yet with a decreased score in terms of $\mathcal{S}^{1}$.
A.16 More Qualitative Segmentations and Statistics for Evaluating the Interpretability of the Concept Detectors
Figures 29, 30, 31, 32, 33, 34, 37, 38, 39 depict qualitative segmentation results obtained using the concept detectors that were learned with the proposed method for the architectures that we studied in this work. Visualizations are obtained using NetworkDissection. We can verify that concept detectors appear to be monosemantic. In some cases, there are more than one concept detector detecting the same concept. However, in most cases, the sets of positively classified samples across detectors are disjoint. Figures 35 and 36 depict qualitative segmentations obtained via NMF and PCA. In some of the depicted examples, it is evident that the identified concepts are less monosemantic. Tables 27, 28, 29 and 30 summarize the statistics obtained by Network Dissection regarding the interpretability of the clusterings considered in this work.
@@ II Method Color Object Part Material Action Scene Texture Total
A.17 More Global Model Explanations via Concept Sensitivity Testing
@@
A.18 More Local Explanations with Concept Contribution Maps
@@
A.19 Details on Counterfactual Explanation Example
For intervention, we make use of (33). To compute the target projected values $t$, we use the training split of Broden. When we perform concept removal (or addition), we choose $t$ to be equal to the lowest (or topmost) $0.005\%$ quantile of projected feature activations on the filter ${\bm{w}}_{i}$. This process ensures that, after the intervention, the signal value in the patch embedding for the concept matches the signal value of samples with or without the concept.
A.20 Details on Toy Model Correction
Network Architecture After each convolutional layer, except the last, there is a Dropout layer with $p=0.3$. All Conv2d layers have a kernel size of 3x3 and a stride of 2, except for the last one, which has a stride of 1. Furthermore, the latent space dimensionality is set to 16 for all convolutional units.
Chess Dataset The total number of images in the dataset is $210$, 67 for bishop, 71 for knight and 72 for rook. The spatial resolution of those images is 85x85. We make a stratified train-test split with the training set ratio set to 0.7.
Network Training We train the network with cross-entropy loss and the Adam optimizer with a learning rate of $0.005$ for $1000$ epochs. In the (poisoned) training set, the model achieves 100% accuracy.
Direction Learning To learn the latent directions, we apply our method by following the learning process described in section A.7. Furthermore, for stable learning, we learn the directions with the Augmented Lagrangian loss scheme of Section 4.5. We optimize $\lambda^{fs}\mathcal{L}^{fs}+\lambda^{cur}\mathcal{L}^{cur}$ with the target constraints $\tau^{ma}=0.8,\tau^{ic}=0.01,\tau^{eac}=0.01,\tau^{mm}=8.0,\tau^{fso}=0.1$, and weights $\lambda^{fs}=2.6$ and $\lambda_{cur}=0.25$.
A.21 Details on Computational Resources
We utilized dedicated computing rigs for our experiments. Experiments for ResNet18 and EfficientNet were run on single-GPU configurations: one NVIDIA A10G (24 GB) and one NVIDIA RTX 4090 (24 GB), respectively. For InceptionV3, VGG16, and ResNet50, we employed a compute rig equipped with four NVIDIA A10G GPUs (4 × 24 GB). The runtime for each of the key processing steps (a), (b), and (d) detailed in Section A.7 varied by architecture: ResNet18 completed in under 2 hours; VGG16 in under 1 hour; ResNet50 in under 10 hours; InceptionV3 in under 13 hours; and EfficientNet in under 5 hours.
A.22 Relation to Sparse Auto-Encoders
The superposition hypothesis Superposition assumes that neural networks linearly represent more features than there are neurons in their hidden layers. In this line of work, a feature is defined to be an abstract property of the input, which may or may not align with human intuition, and is exploited by the network to make predictions. Sparse AutoEncoders (SAEs) DictionaryNLP have been proposed as a tool to take features out of superposition; that is, given the activations of a network’s intermediate layer, SAEs try to linearly decompose these activations in terms of latent feature components under a sparsity objective.
Let ${\bm{x}}\in{\mathbb{R}}^{D}$ denote the activation of a network in a hidden layer of study and $I\gg D$ denote the number of latent features that the network is assumed to represent. While several SAE variants have been proposed, such as SAE1; SAE2; SAE3; matryoshka-saes, in their baseline form DictionaryNLP they first extract latent feature components $\bm{v}$} by:
@@ A.22 Relation to Sparse Auto-Encoders
Despite the conceptual similarities between our method and SAEs, we also have several differences. First, (44) is conceptually acting as a classifier with positive predictions whenever the extracted signal value exceeds the threshold in ${\bm{b}}_{\text{enc}}$. From one perspective, the same equation acts like centering the signal value around the classification threshold for the presence of the concept in the representation. In SAEs, sparsity is enforced in the units of the latent features $\bm{v}$}, while in our approach, the ReLU activation function is replaced with a sigmoid, and thus sparsity is enforced in the semantic space of concepts. Second, in SAEs, ${\bm{W}}_{\text{dec}}$ are learned via a reconstruction objective, while in our approach, signal vectors are learned under the properties of a probabilistic model constrained on the extracted signal values and the feature activations themselves. Third, we also consider constraining each decoding direction to be perpendicular to the signal vectors of other concepts for exact signal value extraction. Finally, we also consider uncertainty region alignment, which additionally exploits the use of the directions by the model.